One water pump can drain a lake in 9 hours. A newer model can drain the same lake in 6 hours. How long does it take for both water pumps together to drain the lake?

Answer by nerdybill(6948)

One water pump can drain a lake in 9 hours. A newer model can drain the same lake in 6 hours. How long does it take for both water pumps together to drain the lake?
.
Let x = time in hours it takes for both to drain lake
then
x (1/9 + 1/6) = 1
Multiply both sides by 54 (common denominator)
x (6 + 9) = 54
x(15) = 54
x = 54/15
x = 3.6 hours
or
x = 3 hours and 36 minutes